Double your wealth. Kant Miss Company is promising its investors that it will double their money every 5 years. What annual rate is Kant Miss promising? Is this investment a good deal? If you invest $250 now and Kant Miss is able to deliver on its promise, how long will it take your investment to reach $34,000? Using the Rule of 72, what annual rate is Kant Miss promising? (Round to the nearest whole percentage.)
Answer #1
Rule of 72 says about implied rate to double money 72/ No of years to double the money
Rule of 72 says about implied rate to double money 72/ 5
Rule of 72 says about implied rate to double money 14.4 %
34000 =250*(1+14.4%)^N
34000/250 =1.144^N
                                                               136.00 =1.144^N
Log 136 N * log 1.144
2.13353891 N* 0.05842602
N= 2.13353891/0.05842602
N=                                   36.52 Years